Find an equation for the line with the given properties.
perpendicular to the line y=1/9x+2; containing the point (3,-3)

Answer by ewatrrr(24785)


Hi


y = (1/9)x+2 |slope of ⊥ line would be -9 (negative reciprocal of 1/9 )
New Line
y = -9x + b |using ordered pair Pt(3,-3) mto solve for b
-3 = -27 + b
24 = b
y = -9x + 24